# Mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium: An effective-dipole analysis

### v1

@@ III Results and Discussion
To investigate how two oppositely charged particles in a more polarizable medium interact, we have repeated our numerical investigations for scenario 2(a) with a sign inversion of $q_{2}$ and with all other parameters remaining unchanged. The charge ratios considered are therefore (a) $q_{2}/q_{1}=-3$, (b) $q_{2}/q_{1}=-8$, (c) $q_{2}/q_{1}=-16$, and (d) $q_{2}/q_{1}=-80$, respectively. The results are summarized in Fig. 7. As shown in the first column, for decreasing interparticle separation, there is generally an increasing deviation of the electrostatic force $F_{2}$ on sphere 2 from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced repulsion. At sufficiently large absolute values of $q_{2}/q_{1}$ (e.g. $q_{2}/q_{1}=-8$, $q_{2}/q_{1}=-16$ and $q_{2}/q_{1}=-80$), the Coulomb attraction is overtaken by the polarization-induced repulsive force at short interparticle separations such that the overall interparticle interaction becomes counterintuitively repulsive. As shown in the second column, for decreasing interparticle separation, the magnitude of induced charge in each sphere as relative to $q_{1}$, i.e. $\widetilde{q}_{i}/q_{1}$, generally increases. For the four different values of $q_{2}/q_{1}$, we observed qualitative differences in the variation of $\widetilde{q}_{i}/q_{1}$ for decreasing interparticle separation: (a) At $q_{2}/q_{1}=-3$, neither the value of $\widetilde{q}_{1}/q_{1}$ nor that of $\widetilde{q}_{2}/q_{1}$ exceeds unity, and the interparticle interaction remains attractive all the way; (b) At $q_{2}/q_{1}=-8$, the magnitudes of induced charge in both spheres are still all the way smaller than $q_{1}$, yet the value of $\widetilde{q}_{1}/q_{1}$ is not far from unity at short interparticle separations. This corresponds to a sufficiently strong effect of polarization for the occurrence of opposite-charge repulsion; (c) At $q_{2}/q_{1}=-16$, the value of $\widetilde{q}_{1}/q_{1}$ exceeds unity at small interparticle separations but the value of $\widetilde{q}_{2}/q_{1}$ is still all the way smaller than unity, and opposite-charge repulsion occurs at short interparticle separations; (d) At $q_{2}/q_{1}=-80$, both the value of $\widetilde{q}_{1}/q_{1}$ and that of $\widetilde{q}_{2}/q_{1}$ exceed unity at certain interparticle separations, and opposite-charge repulsion occurs at short interparticle separations. As shown in the third column, the positions of induced positive charge and induced negative charge in sphere 2 as relative to the sphere’s centre exhibit a non-trivial variation for decreasing interparticle separation, at all four values of $q_{2}/q_{1}$. The results are translated into a pictorial representation in the fourth column, which illustrates the positional order of induced charge and free charge in each sphere. While the positional order of charge in sphere 1 is the same for any interparticle separation or charge ratio, i.e. with the induced positive charge on the left, the free positive charge in the middle, and the induced negative charge on the right, the positional order of charge in sphere 2 exhibits an interesting dependence on the interparticle separation and on the charge ratio, as discussed below:
@@
At (d) $q_{2}/q_{1}=-80$, the polarization of the medium region surrounding sphere 2 is stronger than the three cases described above. For decreasing interparticle separation, the induced positive charge in the medium region between the spheres not only results in an attraction of the induced negative charge $\widetilde{q}_{2}^{-}$ in sphere 2 from the rightmost position to the leftmost position in the corresponding pictorial representation, but also a repulsion of the induced positive charge $\widetilde{q}_{2}^{+}$ in sphere 2 from the leftmost position to the rightmost position. Like the cases of $q_{2}/q_{1}=-8$ and $q_{2}/q_{1}=-16$, this results in a reversal of polarization in sphere 2 and a switch of the electrostatic force from attractive to repulsive, in which case the repulsive interaction between the induced negative charge $\widetilde{q}_{1}^{-}$ in sphere 1 and the induced negative charge $\widetilde{q}_{2}^{-}$ in sphere 2 becomes dominant. At very short interparticle separations, the induced positive charge in the medium region to the right of sphere 2 and the depletion-weakened induced positive charge in the medium region between the spheres together result in a repulsion of the induced positive charge $\widetilde{q}_{2}^{+}$ in sphere 2 from the rightmost position to the middle position in the corresponding pictorial representation.
@@
In general, if the dielectric constant of the surrounding medium is larger than the dielectric constants of two like-charged spheres, i.e. $k_{1}<k_{0}$ and $k_{2}<k_{0}$, the force $F_{2}$ acting on sphere 2 is an amplification of the Coulomb repulsion by a polarization-induced repulsive force. Else if the spheres are oppositely charged, the force $F_{2}$ experienced by sphere 2 is a mitigation of the Coulomb attraction by a polarization-induced repulsive force. It is worth noting that a sign inversion of $q_{2}$ has little impact on the polarization-induced repulsive force, as illustrated in Fig. 6 for the charge ratios $q_{2}/q_{1}=\pm 3$, $q_{2}/q_{1}=\pm 8$, $q_{2}/q_{1}=\pm 16$ and $q_{2}/q_{1}=\pm 80$.
IV Conclusions
Figure 8: Schematic illustration of the four scenarios of electrostatic interactions under consideration: 1(a) $k_{1}>k_{0}$, $k_{2}>k_{0}$ and $q_{2}/q_{1}>0$; 1(b) $k_{1}>k_{0}$, $k_{2}>k_{0}$ and $q_{2}/q_{1}<0$; 2(a) $k_{1}<k_{0}$, $k_{2}<k_{0}$ and $q_{2}/q_{1}>0$; 2(b) $k_{1}<k_{0}$, $k_{2}<k_{0}$ and $q_{2}/q_{1}<0$.
@@ IV Conclusions
By means of numerical simulations and theoretical analysis, we have investigated the mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium, and have analyzed our results in terms of an effective dipole in each particle. Such a dipole depicts how polarization charges of opposite signs are spatially separated. For short inter-particle separations with strong effects of polarization, our findings revealed a significant deviation of the interparticle force from that described by Coulomb’s law and an intriguing correlation of the nature of interparticle interaction with the polarizability of the medium, as summarized in Fig. 8: If the dielectric constant of the surrounding medium is smaller than the dielectric constants of both spheres, i.e. $k_{0}<k_{1}$ and $k_{0}<k_{2}$, the polarization within the system results in an additional attractive interaction that either weakens (or even overcomes) the Coulomb repulsion between two like-charged particles ($q_{2}/q_{1}>0$) or strengthens the Coulomb attraction between two oppositely-charged particles ($q_{2}/q_{1}<0$). Else if the dielectric constant of the surrounding medium exceeds the dielectric constants of both spheres, i.e. $k_{1}<k_{0}$ and $k_{2}<k_{0}$, the polarization within the system results in an additional repulsive interaction that either strengthens the Coulomb repulsion between two like-charged particles ($q_{2}/q_{1}>0$) or weakens (or even overcomes) the Coulomb attraction between two oppositely-charged particles ($q_{2}/q_{1}<0$). In either case, wherever the force on the more polarizing particle switches from repulsive to attractive or vice versa for decreasing inter-particle separation, the effective dipole in this particle undergoes a reversal in its orientation. Our findings provide physical insights into how the electrostatic interaction between two charged dielectric particles is influenced by the spatial distribution of polarization charges in each particle and by the polarizability of the medium, and hold valuable implications for future explorations of multi-body systems.
V Acknowledgements

### v2

@@ IV Mechanisms of interparticle interactions
We have repeated our numerical investigations with a sign inversion of $q_{2}$ and with all other parameters remaining unchanged. The charge ratios considered are therefore (a) $q_{2}/q_{1}=-3$, (b) $q_{2}/q_{1}=-8$, (c) $q_{2}/q_{1}=-16$, and (d) $q_{2}/q_{1}=-80$, respectively. The results, which are summarized in Fig. 9, are analogous to those for scenario B, where in both scenarios there is a crossover between attraction and repulsion. For decreasing interparticle separation, there is generally an increasing deviation of the electrostatic force $F_{2}$ on sphere 2 from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced repulsion. At sufficiently large absolute values of $q_{2}/q_{1}$ (e.g. $q_{2}/q_{1}=-8$, $q_{2}/q_{1}=-16$ and $q_{2}/q_{1}=-80$), the Coulomb attraction is overtaken by the polarization-induced repulsive force at short interparticle separations such that the overall interparticle interaction becomes counterintuitively repulsive.
For decreasing interparticle separation, the magnitude of induced charge in each sphere as relative to $q_{1}$, i.e. $\widetilde{q}_{i}/q_{1}$, generally increases. For the four different values of $q_{2}/q_{1}$, we observed qualitative differences in the variation of $\widetilde{q}_{i}/q_{1}$ for decreasing interparticle separation: (a) At $q_{2}/q_{1}=-3$, neither the value of $\widetilde{q}_{1}/q_{1}$ nor that of $\widetilde{q}_{2}/q_{1}$ exceeds unity, and the interparticle interaction remains attractive all the way; (b) At $q_{2}/q_{1}=-8$, the magnitudes of induced charge in both spheres are still all the way smaller than $q_{1}$, yet the value of $\widetilde{q}_{1}/q_{1}$ is not far from unity at short interparticle separations. This corresponds to a sufficiently strong effect of polarization for the occurrence of opposite-charge repulsion; (c) At $q_{2}/q_{1}=-16$, the value of $\widetilde{q}_{1}/q_{1}$ exceeds unity at small interparticle separations but the value of $\widetilde{q}_{2}/q_{1}$ is still all the way smaller than unity, and opposite-charge repulsion occurs at short interparticle separations; (d) At $q_{2}/q_{1}=-80$, both the value of $\widetilde{q}_{1}/q_{1}$ and that of $\widetilde{q}_{2}/q_{1}$ exceed unity at certain interparticle separations, and opposite-charge repulsion occurs at short interparticle separations.
The positions of induced positive charge and induced negative charge in sphere 2 as relative to the sphere’s centre exhibit a non-trivial variation for decreasing interparticle separation, at all four values of $q_{2}/q_{1}$. While the positional order of charge in sphere 1 is the same for any interparticle separation or charge ratio, i.e. with the induced positive charge on the left, the free positive charge in the middle, and the induced negative charge on the right, the positional order of charge in sphere 2 exhibits an interesting dependence on the interparticle separation and on the charge ratio, as discussed below:
@@
At (d) $q_{2}/q_{1}=-80$, the polarization of the medium region surrounding sphere 2 is stronger than the three cases described above. For decreasing interparticle separation, the induced positive charge in the medium region between the spheres not only results in an attraction of the induced negative charge $\widetilde{q}_{2}^{-}$ in sphere 2 from the rightmost position to the leftmost position in the corresponding pictorial representation, but also a repulsion of the induced positive charge $\widetilde{q}_{2}^{+}$ in sphere 2 from the leftmost position to the rightmost position. Like the cases of $q_{2}/q_{1}=-8$ and $q_{2}/q_{1}=-16$, this results in a reversal of polarization in sphere 2 and a switch of the electrostatic force from attractive to repulsive, in which case the repulsive interaction between the induced negative charge $\widetilde{q}_{1}^{-}$ in sphere 1 and the induced negative charge $\widetilde{q}_{2}^{-}$ in sphere 2 becomes dominant.
At very short interparticle separations, the induced positive charge in the medium region to the right of sphere 2 and the depletion-weakened induced positive charge in the medium region between the spheres together result in a repulsion of the induced positive charge $\widetilde{q}_{2}^{+}$ in sphere 2 from the rightmost position to the middle position in the corresponding pictorial representation.
@@
If the dielectric constant of the surrounding medium is larger than the dielectric constants of two like-charged spheres, i.e. $k_{1}<k_{0}$ and $k_{2}<k_{0}$, the force $F_{2}$ acting on sphere 2 is an amplification of the Coulomb repulsion by a polarization-induced repulsive force. Else if the spheres are oppositely charged, the force $F_{2}$ experienced by sphere 2 is a mitigation of the Coulomb attraction by a polarization-induced repulsive force. As illustrated in Fig. 10, a sign inversion of $q_{2}$ has little impact on the polarization-induced repulsive force.
V Conclusions
@@
By means of numerical simulations and theoretical analysis, we have investigated the mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium. For short inter-particle separations with strong effects of polarization, our findings revealed a significant deviation of the interparticle force from that described by Coulomb’s law and an intriguing correlation of the nature of interparticle interaction with the polarizability of the medium, as summarized in Fig. 4. Wherever the force on the more polarizing particle switches from repulsive to attractive or vice versa for decreasing interparticle separation, the effective dipole in this particle undergoes a reversal in its orientation. Our findings, which have also revealed possible depletion effects of polarization charge in the medium for near-touching spheres, contribute to a comprehensive understanding of the mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium, and have valuable implications for future explorations of multi-body systems.
For a pair of polarizable ions in a vacuum, the impossibility of opposite-charge repulsion has been derived analytically based on the empirical polarizability-volume relation of a polarizable-ion model Chan (2020). As possible future work, for a system of two charged dielectric spheres, it might be possible to derive the impossibility of opposite-charge repulsion in a comparatively less polarizable medium and of like-charge attraction in a comparatively more polarizable medium if we are able to obtain an empirical relation between the effective-dipole moment and local electric field for each sphere.
VI Conflicts of interest
There are no conflicts of interest to declare.
VII Acknowledgements